# Totally geodesic submanifolds of the homogeneous nearly Kähler 6-manifolds and their $\mathsf{G}_{2}$-cones

### v2

@@ 7.1. The complex projective space
If $X\in\mathfrak{v}$ is a unit vertical vector, we may assume by means of the isotropy representation that $X=e_{1}$. Since the spectrum of the Jacobi operator $R_{e_{1}}\colon\mathfrak{p}\ominus\mathbb{R}e_{1}\to\mathfrak{p}\ominus\mathbb{R}e_{1}$ consists of the eigenvalues $2$, with eigenspace $\mathbb{R}e_{2}$, and $1/8$, with eigenspace $\mathfrak{p}_{2}$, the claim follows from Remark 3.3. Similarly, if $X$ is horizontal we may suppose that $X=e_{3}$, and in this case the eigenvalues of $R_{e_{3}}\colon\mathfrak{p}\ominus\mathbb{R}e_{3}\to\mathfrak{p}\ominus\mathbb{R}e_{3}$ are $\frac{1}{8}$, with eigenspace $\mathfrak{p}_{1}$, $1$, with eigenspace $\mathbb{R}e_{4}$, and $5/8$, with eigenspace $\operatorname{span}\{e_{5},e_{6}\}$, so the result also holds in this case. ∎
@@
Assume that $\mathfrak{v}\subseteq\mathfrak{p}_{2}$. Using the isotropy representation if necessary, we can assume that $e_{3}\in\mathfrak{v}$. One sees that the kernel of the Cartan operator $C_{X}$ is spanned by $e_{3}$ and $e_{4}$, so we must have $\mathfrak{v}=\operatorname{span}\{e_{3},e_{4}\}=\mathfrak{p}_{\mathsf{SU}(2)\cdot o}$, since $\Sigma$ has constant curvature, which means that $\Sigma=\mathsf{SU}(2)\cdot o$.
@@ 7.2. The flag manifold
First, assume that $\lambda>0$, so the three eigenvalues given above are pairwise distinct. We will prove that $\mathfrak{v}$ coincides with the kernel of the Cartan operator $C_{X}$.
@@
Now, suppose that $\lambda=0$, and we deduce that $e_{1}\in\mathfrak{v}$. Since the intersection of $\mathfrak{v}$ with $\mathfrak{p}_{2}\oplus\mathfrak{p}_{3}$ is at least two-dimensional, we can use the isotropy representation to assume that there is a tangent vector of the form $Y=e_{3}+\mu e_{5}$, where $\mu\in\mathbb{R}$. However, $\mu=0$, for if $\mu\neq 0$, by means of the full isotropy representation we can conjugate $\mathfrak{v}$ to a new totally geodesic subspace containing a tangent vector of the form $e_{1}+\mu e_{3}$, and we may use the previous case to derive a contradiction. Now this yields that $e_{1}+e_{3}$ is also in $\mathfrak{v}$, which is yet another contradiction. We conclude that this case is also not possible, and therefore there are no codimension two totally geodesic subspaces in $\mathfrak{p}$. ∎
@@
Let $\Sigma\subseteq\mathsf{F}(\mathbb{C}^{3})$ be a complete totally geodesic submanifold of dimension three. Then $\Sigma$ is congruent to either $\mathsf{F}(\mathbb{R}^{3})$ or to $\mathsf{S}^{3}_{\mathbb{C},1/4}(\sqrt{2})$.
@@
First, assume that $\lambda>0$, so the three eigenvalues given above are pairwise distinct. We will prove in this case that either $\Sigma=\mathsf{SU}(2)_{(1,0,1)}\cdot o$ or $\mathfrak{v}\subseteq\ker C_{X}$. Indeed, if there is a vector of the form $Y=\varepsilon e_{2}-\lambda\varepsilon e_{4}+\sqrt{1+\lambda^{2}}e_{6}$ in $\mathfrak{v}$, where $\varepsilon\in\{\pm 1\}$, then we may construct a basis of $\mathfrak{v}$ by adding the vector
@@
Let $\mathfrak{v}$ be a totally geodesic subspace of dimension $2$ in $\mathfrak{p}$. We define $r\in\{1,2,3\}$ to be the largest number such that there exists a vector $X\in\mathfrak{v}$ that projects nontrivially in $r$ of the irreducible submodules of $\mathfrak{p}$.
The case $r=3$. Let $X\in\mathfrak{v}$ be a vector that projects nontrivially on each of the submodules $\mathfrak{p}_{i}$. Then, by means of the isotropy representation, we know that $X$ is (up to $\mathsf{T}^{2}$-conjugacy and scaling) of the form $X=e_{1}+a_{3}e_{3}+a_{5}e_{5}+a_{6}e_{6}$, where $a_{3}$, $a_{5}^{2}+a_{6}^{2}\neq 0$.
@@
The case $r=2$. We can find a vector $X\in\mathfrak{v}$ that projects nontrivially on two of the three irreducible submodules. By using the full isotropy representation and rescaling, we can assume that $X=e_{1}+\lambda e_{3}$ for a certain $\lambda>0$. Take any vector $Y\in\mathfrak{v}\ominus\mathbb{R}X$.
@@ 7.3. The almost product $\mathsf{S}^{3}\times\mathsf{S}^{3}$
Suppose that $\mathfrak{v}$ contains a nonzero vector $X\in\mathfrak{p}_{1}$ (respectively, $X\in\mathfrak{p}_{2}$). Since the isotropy representation of $\Delta\mathsf{SU}(2)$ is transitive in the unit sphere of $\mathfrak{p}_{1}$ (respectively, $\mathfrak{p}_{2}$), we may suppose that $X=e_{1}$ (respectively, $X=e_{4}$). Note that the matrices of $R_{e_{1}}$ and $R_{e_{4}}$ are given by $R_{e_{1}}=\operatorname{diag}(0,3/4,3/4,0,1/12,1/12)$ and $R_{e_{4}}=\operatorname{diag}(0,1/12,1/12,0,3/4,3/4)$, which means that $\mathfrak{v}=(\mathfrak{v}\cap\mathbb{R}e_{1})\oplus(\mathfrak{v}\cap\operatorname{span}\{e_{2},e_{3}\})\oplus(\mathfrak{v}\cap\mathbb{R}e_{4})\oplus(\mathfrak{v}\cap\operatorname{span}\{e_{5},e_{6}\})$ in both cases. This last equation implies directly that $\mathfrak{v}$ is well-positioned. ∎
@@
One sees that the product $fg$ is precisely the characteristic polynomial of $R_{X}\in\mathsf{End}(\mathfrak{p}\ominus\mathbb{R}X)$, so in particular we obtain that $f(R_{X})g(R_{X})=0$. Furthemore, $f$ and $g$ are relatively prime. As a consequence, because $\mathfrak{v}$ is an $R_{X}$-invariant subspace, it may be decomposed as $\mathfrak{v}=\mathbb{R}X\oplus(\mathfrak{v}\cap\ker f(R_{X}))\oplus(\mathfrak{v}\cap\ker g(R_{X}))$. It turns out that
@@
We now let $Y=e_{1}+\lambda e_{4}\in\mathfrak{v}$, where $\lambda\in\mathbb{R}\setminus\{0\}$. The Jacobi operator $R_{Y}\in\mathsf{End}(\mathfrak{p}\ominus\mathbb{R}Y)$ has three different eigenvalues 00, $\frac{1+\lambda^{2}}{12}$ and $\frac{3(1+\lambda^{2})}{4}$, with corresponding eigenspaces $\mathbb{R}(\lambda e_{4}-e_{1})$, $\operatorname{span}\{\lambda e_{2}-e_{5},\lambda e_{3}-e_{6}\}$ and $\operatorname{span}\{e_{2}+\lambda e_{5},e_{3}+\lambda e_{6}\}$. Note that the isotropy subgroup $\mathsf{K}_{e_{1}}$ fixes $Y$ and acts transitively on the set of lines in both $\operatorname{span}\{\lambda e_{2}-e_{5},\lambda e_{3}-e_{6}\}$ and $\operatorname{span}\{e_{2}+\lambda e_{5},e_{3}+\lambda e_{6}\}$. Now, the fact that $\dim\mathfrak{v}=3$ implies that $\mathfrak{v}$ must contain a nonzero eigenvector from one of the last two eigenspaces. Our next step is to prove that $\lambda^{2}=3$ or $\lambda^{2}=1/3$.
If $\mathfrak{v}\cap\operatorname{span}\{\lambda e_{2}-e_{5},\lambda e_{3}-e_{6}\}\neq 0$, we can conjugate by an element of $\mathsf{K}_{e_{1}}$ to assume that $Z=\lambda e_{2}-e_{5}\in\mathfrak{v}$ is tangent to $\Sigma$. If we assume that $\lambda^{2}\notin\left\{3,1/3\right\}$, then we can give a basis of $\mathfrak{v}$ with the vectors
@@
a contradiction, so we deduce that $\lambda^{2}\in\left\{3,1/3\right\}$.
@@
Let $\Sigma$ be a complete totally geodesic submanifold of the $\mathsf{G}_{2}$-cone $\widehat{M}$. By Corollary 6.10, $\Sigma$ is either a hypersurface of $\widehat{M}$ or the cone of a maximal totally geodesic submanifold of $M$. The first case is not possible due to [32, Theorem 1.2], so we conclude that $\Sigma=\widehat{S}$ for a maximal totally geodesic submanifold $S\subseteq M$, which yields the desired result. ∎

### v3

@@ 7.1. The complex projective space
If $X\in\mathfrak{v}$ is a unit vertical vector, we may assume by means of the isotropy representation that $X=e_{1}$. Since the spectrum of the Jacobi operator $R_{e_{1}}\colon\mathfrak{p}\ominus\mathbb{R}e_{1}\to\mathfrak{p}\ominus\mathbb{R}e_{1}$ consists of the eigenvalues $2$, with eigenspace $\mathbb{R}e_{2}$, and $1/8$, with eigenspace $\mathfrak{p}_{2}$, the claim follows from Remark 3.3. Similarly, if $X$ is horizontal we may suppose that $X=e_{3}$, and in this case the eigenvalues of $R_{e_{3}}\colon\mathfrak{p}\ominus\mathbb{R}e_{3}\to\mathfrak{p}\ominus\mathbb{R}e_{3}$ are $1/8$, with eigenspace $\mathfrak{p}_{1}$, $1$, with eigenspace $\mathbb{R}e_{4}$, and $5/8$, with eigenspace $\operatorname{span}\{e_{5},e_{6}\}$, so the result also holds in this case. ∎
@@
Assume that $\mathfrak{v}\subseteq\mathfrak{p}_{2}$. Using the isotropy representation if necessary, we can suppose that $e_{3}\in\mathfrak{v}$. One sees that the kernel of the Cartan operator $C_{X}$ is spanned by $e_{3}$ and $e_{4}$, so we must have $\mathfrak{v}=\operatorname{span}\{e_{3},e_{4}\}=\mathfrak{p}_{\mathsf{SU}(2)\cdot o}$, since $\Sigma$ has constant curvature, which means that $\Sigma=\mathsf{SU}(2)\cdot o$.
@@ 7.2. The flag manifold
First, assume that $\lambda>0$, so the three eigenvalues given above are pairwise distinct. We prove that $\mathfrak{v}$ coincides with the kernel of the Cartan operator $C_{X}$.
@@
Now, suppose that $\lambda=0$, which implies that $e_{1}\in\mathfrak{v}$. Since the intersection of $\mathfrak{v}$ with $\mathfrak{p}_{2}\oplus\mathfrak{p}_{3}$ is at least two-dimensional, we can use the isotropy representation to assume that there is a tangent vector of the form $Y=e_{3}+\mu e_{5}$, where $\mu\in\mathbb{R}$. However, $\mu=0$, for if $\mu\neq 0$, by means of the full isotropy representation we can conjugate $\mathfrak{v}$ to a new totally geodesic subspace containing a tangent vector of the form $e_{1}+\mu e_{3}$, and we may use the previous case to derive a contradiction. Now this yields that $e_{1}+e_{3}$ is also in $\mathfrak{v}$, which is yet another contradiction. We conclude that this case is also not possible, and therefore there are no codimension two totally geodesic subspaces in $\mathfrak{p}$. ∎
@@
Let $\Sigma$ be a complete three-dimensional totally geodesic submanifold of the flag manifold $\mathsf{F}(\mathbb{C}^{3})$. Then $\Sigma$ is congruent to either $\mathsf{F}(\mathbb{R}^{3})$ or to $\mathsf{S}^{3}_{\mathbb{C},1/4}(\sqrt{2})$.
@@
First, assume that $\lambda>0$, so the three eigenvalues given above are pairwise distinct. We prove in this case that either $\Sigma=\mathsf{SU}(2)_{(1,0,1)}\cdot o$ or $\mathfrak{v}\subseteq\ker C_{X}$. Indeed, if there is a vector of the form $Y=\varepsilon e_{2}-\lambda\varepsilon e_{4}+\sqrt{1+\lambda^{2}}e_{6}$ in $\mathfrak{v}$, where $\varepsilon\in\{\pm 1\}$, then we may construct a basis of $\mathfrak{v}$ by adding the vector
@@
Let $\mathfrak{v}$ be a totally geodesic subspace of dimension $2$ in $\mathfrak{p}$. We define $r\in\{1,2,3\}$ to be the largest number such that there exists a vector $X\in\mathfrak{v}$ that has nontrivial projection onto $r$ of the irreducible submodules of $\mathfrak{p}$.
The case $r=3$. Let $X\in\mathfrak{v}$ be a vector that projects nontrivially onto each of the submodules $\mathfrak{p}_{i}$. Then, by means of the isotropy representation, we know that $X$ is (up to $\mathsf{T}^{2}$-conjugacy and scaling) of the form $X=e_{1}+a_{3}e_{3}+a_{5}e_{5}+a_{6}e_{6}$, where $a_{3}$, $a_{5}^{2}+a_{6}^{2}\neq 0$.
@@
The case $r=2$. We can find a vector $X\in\mathfrak{v}$ that projects nontrivially onto two of the three irreducible submodules. By using the full isotropy representation and rescaling, we can assume that $X=e_{1}+\lambda e_{3}$ for a certain $\lambda>0$. Take any vector $Y\in\mathfrak{v}\ominus\mathbb{R}X$.
@@ 7.3. The almost product $\mathsf{S}^{3}\times\mathsf{S}^{3}$
Suppose that $\mathfrak{v}$ contains a nonzero vector $X\in\mathfrak{p}_{1}$ (respectively, $X\in\mathfrak{p}_{2}$). Since the isotropy representation of $\Delta\mathsf{SU}(2)$ is transitive on the unit sphere of $\mathfrak{p}_{1}$ (respectively, $\mathfrak{p}_{2}$), we may suppose that $X=e_{1}$ (respectively, $X=e_{4}$). Note that the matrices of $R_{e_{1}}$ and $R_{e_{4}}$ are given by $R_{e_{1}}=\operatorname{diag}(0,3/4,3/4,0,1/12,1/12)$ and $R_{e_{4}}=\operatorname{diag}(0,1/12,1/12,0,3/4,3/4)$, which means that $\mathfrak{v}=(\mathfrak{v}\cap\mathbb{R}e_{1})\oplus(\mathfrak{v}\cap\operatorname{span}\{e_{2},e_{3}\})\oplus(\mathfrak{v}\cap\mathbb{R}e_{4})\oplus(\mathfrak{v}\cap\operatorname{span}\{e_{5},e_{6}\})$ in both cases. This last equation implies directly that $\mathfrak{v}$ is well-positioned. ∎
@@
One sees that the product $fg$ is precisely the characteristic polynomial of $R_{X}\in\operatorname{End}(\mathfrak{p}\ominus\mathbb{R}X)$, so in particular we obtain that $f(R_{X})g(R_{X})=0$. Furthermore, $f$ and $g$ are relatively prime. As a consequence, because $\mathfrak{v}$ is an $R_{X}$-invariant subspace, it may be decomposed as $\mathfrak{v}=\mathbb{R}X\oplus(\mathfrak{v}\cap\ker f(R_{X}))\oplus(\mathfrak{v}\cap\ker g(R_{X}))$. It turns out that
@@
We now let $Y=e_{1}+\lambda e_{4}\in\mathfrak{v}$, where $\lambda\in\mathbb{R}\setminus\{0\}$. The Jacobi operator $R_{Y}\in\operatorname{End}(\mathfrak{p}\ominus\mathbb{R}Y)$ has three different eigenvalues 00, $\frac{1+\lambda^{2}}{12}$ and $\frac{3(1+\lambda^{2})}{4}$, with corresponding eigenspaces $\mathbb{R}(\lambda e_{4}-e_{1})$, $\operatorname{span}\{\lambda e_{2}-e_{5},\lambda e_{3}-e_{6}\}$ and $\operatorname{span}\{e_{2}+\lambda e_{5},e_{3}+\lambda e_{6}\}$. Note that the isotropy subgroup $\mathsf{K}_{e_{1}}$ fixes $Y$ and acts transitively on the set of lines in both $\operatorname{span}\{\lambda e_{2}-e_{5},\lambda e_{3}-e_{6}\}$ and $\operatorname{span}\{e_{2}+\lambda e_{5},e_{3}+\lambda e_{6}\}$. Now, the fact that $\dim\mathfrak{v}=3$ implies that $\mathfrak{v}$ must contain a nonzero eigenvector from one of the last two eigenspaces. Our next step is to prove that $\lambda^{2}=3$ or $\lambda^{2}=1/3$.
If $\mathfrak{v}\cap\operatorname{span}\{\lambda e_{2}-e_{5},\lambda e_{3}-e_{6}\}\neq 0$, we can conjugate by an element of $\mathsf{K}_{e_{1}}$ to assume that $Z=\lambda e_{2}-e_{5}\in\mathfrak{v}$ is tangent to $\Sigma$. If we suppose that $\lambda^{2}\notin\left\{3,1/3\right\}$, then we can give a basis of $\mathfrak{v}$ with the vectors
@@
a contradiction, so we obtain that $\lambda^{2}\in\left\{3,1/3\right\}$.
@@
Let $\Sigma$ be a complete totally geodesic submanifold of the $\mathsf{G}_{2}$-cone $\widehat{M}$. By Corollary 6.10, $\Sigma$ is either a hypersurface of $\widehat{M}$ or the cone of a maximal totally geodesic submanifold of $M$. The first case is not possible due to [31, Theorem 1.2], so we conclude that $\Sigma=\widehat{S}$ for a maximal totally geodesic submanifold $S\subseteq M$, which yields the desired result. ∎